OpenStudy (anonymous):

0.5e^(0.08x)=40 how do you solve

OpenStudy (mertsj):

Take the natural log of both sides.

OpenStudy (vincent-lyon.fr):

...after multiplying by 2.

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

0.5e^(0.08x)=40 divide each side by 0.5 e^(0.08x)=80 take ln of both sides to get rid of e 0.08x=ln(80) divide both sides y 0.08 (when you plug this into your calc. do the ln(80) first then divide that answer by 0.08 to find x)